(2) Essam is looking to purchase an apartment for a price of 125,000. Essam plans to put down €20,000 and will pay the rest by taking on a 12% percent two-year bank loan, what is the monthly payment on this car loan?
Answer #1

amount financed

125000-20000

105000

amount of monthly payment

Using rate function in MS excel

pmt(rate,nper,pv,fv,type) rate = 12/12 =1% nper =12*2 =24 pv = 105000 fv =0 type =0

PMT(1%,24,105000,0,0)

($4,942.71)

Monthly payment

($4,942.71)